Suppose the spot rate of Polish Zloty to the US dollar is 4.0 and the forward rate in the 180 day market of the Polish Zloty is 3.5 per dollar. The Polish Zloty is then selling at a _____________ to the spot rate. Express your answer as a decimal not a percentage.
Answer #1

Forward Discount = [Spot Rate - Forward Rate] / Spot Rate

= [4 - 3.5] / 4 = 0.5 / 4 = 0.125, or 12.50%

So, the polish zloty is selling to a discount of 0.125 or 12.5% to the spot rate.